Grade 12th passTrigonometry

Simplify these expressions and find their values if x = 3, a = – 1, b = – 2.
2 – 8x + 4x + 4

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

From the question it is given that x = 3
We have,
= 2 + 4–8x + 4x
= 6–4x
Then, substitute the value of x in the equation
= 6 - (4 × 3)
= 6 - 12
= -6

Thanks